OpenStudy (anonymous):

given f(x)=-3x+2 and g(x)=2x+9 find (g*f)(x)

OpenStudy (anonymous):

(g*f)(x) = g(x)*f(x) (-3x+2)*(2x+9) =(-6x^2 -23x+18)